OpenStudy (anonymous):

Show your work and write a short explanation with each calculation. Part I: Calculate the energy change (q) of the surroundings (water) using the enthalpy equation qwater = m × c × ΔT. We can assume that the specific heat capacity of water is 4.18 J / (g × °C) and the density of water is 1.00 g/mL. The water has absorbed the heat of the metal. So, qwater = qmetal Using the formula qmetal = m × c × ΔT, calculate the specific heat of the metal. Use the data from your experiment for the metal in your calculation.

OpenStudy (anonymous):

Part II: Calculate the energy change (q) of the surroundings (water) using the enthalpy equation qwater = m × c × ΔT. We can assume that the specific heat capacity of water is 4.18 J / (g × °C) and the density of water is 1.00 g/mL. Show ALL your work. The water has absorbed the heat of the metal. So, qwater = qunknown metal Using the formula qunknown metal = m × c × ΔT, calculate the specific heat of the metal. Use the data from your experiment for the unknown metal in your calculation. Show ALL your work.

OpenStudy (anonymous):

I need an answer literally ASAP.

OpenStudy (cuanchi):

we need your data about mass of water, change of temperature, and mass of metal to be able to calculate the C of the metal

OpenStudy (anonymous):

Mass of water would be for the KNOWN metal 25mL; Mass of the Metal would be 41.664g; and the temperature of the metal would be 100 degrees C and the temperature of the mixture would be 34.5 degrees C. I wasn't sure which of the last two that you needed.

OpenStudy (anonymous):

For part 2: UNKNOWN METAL The mass of the metal is 25.605 The temperature of the mixture is 29.1 degrees C The temperature of the metal is 100 degrees C The Mass of water would be 25mL

OpenStudy (cuanchi):

qwater = qmetal q= m C ΔT \[q=m _{1}C _{1}\Delta T=m _{2}C _{2}\Delta T\] m1 and C1 are the water mass and heat capacity m2 and C2 are the metal mass and heat capacity the ΔT is the same in both side of the equation \[m _{1}C _{1}=m _{2}C _{2}\] \[C _{2}=\frac{ m _{1}C _{1} }{m _{2} }\] \[C _{2}=\frac{ 25 g 4.18J/C }{41.664 g }\]= 2.51 J/C

OpenStudy (anonymous):

Thanks you.